# Domain Transfer Through Image-to-Image Translation for Uncertainty-Aware Prostate Cancer Classification

### v1

@@ 4.2.2 Evidential Focal Loss
Patch-driven filtering: Given the uncertainty for each training patch, we simply eliminate $x\%,x\in[10,20]$ of the patches with highest uncertainty and retrain the model on the rest of the samples on the training set.
@@ 5 Experiments & Setup
Data Split: For the domain transfer task, as mentioned in Section 3.2 and 4, the resampled T2 and ADC “images" with size $256\times 256$ from both ProstateX and local datasets is used for training the ACL-GAN model. Particularly, we allocate 90% of images in both datasets as training and keep 10% as validation set to avoid overfitting the ACL-GAN model. Importantly, we ensure that the images corresponding to each patient are exclusively present in either the training or validation set, but not both. To improve the robustness and enhance the ability of the ACL-GAN to capture feature-level representations of 1.5T images, we use all data from our local hospital. However, it is important to note that this approach does not yield any additional impacts on the subsequent classification task. The model only modifies image regions that have visual differences caused by acquisition parameters of various MRI machines, but it does not alter the context of the prostate itself. For the PCa classification task, we use cropped and augmented T2 and ADC “patches" from both datasets. As mentioned before, this includes 204 ProstateX patients translated to 1.5T, as well as 104 patients from our local hospital captured in 1.5T. Regarding the data split for the classification, we keep patches of 34 patients from our local center as test set. From the remaining patches (70 local patients and all ProstateX patients, we allocate 80% for training and 20% for validation, assuring patches from the same patient not included in both of these sets.
Domain Transfer: The first experiment is translating the ProstateX MRI data from 3.0T to 1.5T using our proposed ACL-GAN model through the conversion process mentioned previously. We then evaluate the effectiveness of this approach by using the translated MRI data in a downstream binary classification task for clinically significant PCa.
Classification: We divide our classification experiments into two categories. In the first category, we use the conventional training paradigm without any filtering or uncertainty estimation. We use different model architectures for these experiments as discussed in Section 4.2.1. In the second category, we use the dataset filtering method and evidential focal loss proposed in Section 4.2.2 for training our models. We select two methods (MpMRI and M.S. MpMRI) that achieve the best performance from the first category to conduct experiments in the second category. Additionally, we conduct several ablation studies and report the results. Finally, we focus on dataset filtering during deployment and examine how this technique affects the classification performance of the test data.
@@ 5.1 Experimental Details
We train two ACL-GAN models separately for T2 and ADC images as part of our domain transfer framework. The optimizer used for both models is Stochastic Gradient Descent with Adam update rule (Kingma and Ba, 2014), with an initial learning rate of 0.0001 and weight decay of 0.0001 to prevent overfitting. The batch size is 3 and are trained for 30,000 epochs. Moreover, when training the model for T2 images, we set the $\lambda_{mask}=0.0025,\lambda_{idt}=1,\lambda_{acl}=0.2$ in Equation (1) and lower and upper mask threshold to be 0.005 and 0.1, respectively. When training the model for ADC images, the value of $\lambda_{mask},\lambda_{idt},\lambda_{acl}$ are the same as in the T2 model with lower and upper mask threshold is set to 0.001 and 0.005, respectively. The Least-Square (LS) loss (Mao et al., 2017) is utilized to calculate $\mathcal{L}_{adv}$ and $\mathcal{L}_{acl}$ in Equation (1).
Converting 3.0T to 1.5T: Once we have obtained two ACL-GAN models, we need to standardize the acquisition parameters of 3.0T prostate MRIs to match those of the 1.5T data in our local dataset. To achieve this, we divide the original MRI into multiple 2D grayscale slices. For each 2D slice, we use the generator $G_{T}$ and a noise vector $z$ randomly sampled from $\mathcal{N}(0,1)$ to translate the slice to 1.5T, e.g., $I_{\text{1.5T}}=G_{T}(I_{\text{3.0T}},z)$ in Section 4.1. We repeat this process for all 2D slices and then stack them back together to reconstruct the 3D MRI for each patient. The voxel spacing remains unchanged before and after the translation process. The above process is applied to both T2 and ADC data.
All classification models are trained with Stochastic Gradient Descent with Adam, and batch normalization is used to speed up convergence. In the first category of classification experiments, the traditional focal loss (Lin et al., 2017) with $\gamma=2$ is used to train the model. Specifically, all models except co-teaching are trained for 300 epochs with a learning rate of 0.0001, weight decay of 0.01, and a batch size of 10. To train the co-teaching model, we set the noise rate to 0.1; the forget rate $\tau=0.1$; the number of epochs for linear drop rate $T_{k}=10$ in Equation (2). The model is trained for 300 epochs; the batch size is set to 10, and the learning rate is 0.00001.
For experiments in the second category, to train the “MpMRI” model for patch-driven filtering, we set the learning rate to 0.0001; weight decay to 0.01; total training epochs to 300, and batch size to 10. The class weights $\boldsymbol{\beta}$ in Equation (6) are set to [0.25, 0.75] for filtering 10%, and [0.25, 1.25] for filtering 20% of the training data. For patient-driven filtering, all parameters are the same except the class weights $\boldsymbol{\beta}$ are set to [0.25, 1] for filtering 10% of the training data. Last but not least, we set the initial learning rate to 0.0001; total training epochs to 300; batch size to 10; the learning rate decayed by a factor of 0.1 for every 200 epochs, and the class weights $\boldsymbol{\beta}$ are set to [0.25, 1] for filtering 20% of the training data.
@@ 5.2 Evaluation
The traditional classification metrics e.g., accuracy, sensitivity, specificity, and AUC for this task. Reporting the patient-level performance are more relevant to the real clinical setting. However, since we used patches as input to the model, to be able to calculate the performance metrics at patient-level, we need to aggregate individual results obtained from the individual patches. To achieve this, we first use the classifier to predict test patches, which are then grouped sequentially into $x$ groups, for each patient (20 probabilities due to the augmentation mentioned in Section 3.2) and then compute the median probability $\tilde{p}_{i}$ as the aggregated probability of that patient. Finally, a threshold of 0.5 is used to determine whether the patient has PCa or not, i.e., assigning label 1 if $\tilde{p}_{i}>0.5$, and 0 otherwise.
@@ 6 Results and Discussion
In this section, we report patient-based classification results; the performance of our methods on patches are reported in Appendix A.
6.1 Translated samples from Domain Transfer
Figure 5 shows the visualization of the difference between original 3.0T T2 and ADC images in the ProstateX Challenge and the corresponding translated 1.5T T2 and ADC images using our proposed domain transfer framework. As shown, domain transfer reduces the image contrast and results in loss of minor details in the original 3.0T images. However, to evaluate the effectiveness of the domain transfer framework, a user study involving radiologists’ interpretation is necessary. Our proposed domain transfer framework shows potential improvements over the baseline model SDNet for the following reason. In SDNet (Grebenisan et al., 2021), modality features from randomly selected 1.5T images are merged with anatomical features from 3.0T rather than from the whole distribution of all 1.5T images. In contrast, our method learns the overall data distribution in an adversarial manner, capturing the entire distribution of 1.5T images and performing the translation. Moreover, our method ensures the translated image contains the crucial features of the original image, and the generator only modifies certain parts of the image. Our image translation method is thus more suitable for this task and can be further validated by the classification performance presented in the following sections.
((a)) Original T2 image in 3.0T
((b)) Translated T2 image in 1.5T
((c)) Original ADC image in 3.0T
((d)) Translated ADC image in 1.5T
Figure 5: Translation from 3.0T T2 and ADC images to 1.5T-like images of a random patient in the “ProstateX” Challenge dataset. left two figures: the original 3.0T T2 image, and the same image that is translated to 1.5T. right two figures: the original 3.0T ADC image, and the same image that is translated to 1.5T.
@@ 6.2 PCa classification without filtering
Table 1 summarizes the main results of this study, which contains the PCa classification performance of all experiments we conducted. The table is divided into three sections. The first section corresponds to experiments conducted without using either the dataset filtering or uncertainty estimation, as described in the first category of Section 5. The second and third sections represent experiments with dataset filtering and uncertainty estimation described in the second category of Section 5. In the first section of Table 1, we observe that the AUC of using the co-teaching framework with “MpMRI” architecture as the base model achieves the best AUC and outperforms the baseline.
We also noticed that the sensitivity increases by approximately 50% while the specificity only decreases  10% for our co-teaching model compared to the baseline model, indicating our model has better learning abilities for classifying both positive and negative data samples. In the training process, we adopt a greedy approach of assuming 10% of the samples to be noisy. Consequently, both models need to designate a portion of the data in each batch as “clean" to update the parameters. This strategy allows our model to prioritize learning from the clean data, leading to enhanced robustness.
Ablation Study: We embed the results of the ablation study in the first section of Table 1. Our ablation here is two-fold: alteration of the number of input modalities, and alteration of the architecture of the model. To examine the effect of data modalities on the classification performance, we compare the T2-only model with (Vol.)MpMRI and M.S. MpMRI models, both use T2 and ADC patches as input. We observe a significant improvement in classification performance with the addition of the ADC modality, suggesting that multi-modal information is useful in guiding the model to classify clinically significant PCa. To examine the effect of model architecture on classification performance, we compare the MpMRI and M.S. MpMRI models, which have different architectures, and the co-teaching model. We find that the model with simpler inputs, MpMRI, performs better, and the results can be further improved by using the co-teaching framework.
Data F.R. F.M. Acc. Sen. Spec. AUC
SDNet(baseline) T2 0% N/A 79.4 28.6 92.6 76.2$\pm$17.5
T2-only T2 0% N/A 64.7 71.4 63.0 77.8$\pm$18.0
MpMRI T2+ADC 0% N/A 79.4 85.7 77.8 84.7$\pm$15.5
Vol. MpMRI T2+ADC 0% N/A 67.6 71.4 66.7 68.9$\pm$26.1
M.S. MpMRI T2+ADC 0% N/A 73.5 71.4 74.1 82.5$\pm$14.3
MpMRI+co-teaching T2+ADC 0% N/A 82.3 85.7 81.2 88.4$\pm$10.6
MpMRI T2+ADC 10% patch 82.4 85.7 81.5 85.7$\pm$9.9
M.S.MpMRI T2+ADC 10% patch 82.4 71.4 85.2 83.6$\pm$13.5
MpMRI T2+ADC 20% patch 85.3 100 81.5 98.4$\pm$1.6
M.S. MpMRI T2+ADC 20% patch 85.3 71.4 88.9 92.6$\pm$7.4
MpMRI T2+ADC 10% patient 88.2 100 85.2 92.6$\pm$7.4
M.S. MpMRI T2+ADC 10% patient 85.3 100 81.5 86.2$\pm$9.0
MpMRI T2+ADC 20% patient 73.5 85.7 70.4 86.8$\pm$12.6
M.S. MpMRI T2+ADC 20% patient 73.5 71.4 74.1 84.6$\pm$14.2
Table 1: Patient-based results of all performed experiments. Acc., Sen., Spec., and AUC are the shorts for Accuracy, Sensitivity, Specificity, and AUC, respectively. Standard deviations are computed from the 95% bootstrap confidence interval with $n=3000$ samples. The “F.R.” and “F.M.” represents the filtering rate and the filtering method, respectively. The best results for each section are in black bold; blue bold, and red bold separately. All units of the numeric values are in %.
@@ 6.3 PCa classification with filtration
In this section, we conduct experiments using two different architectures (MpMRI and M.S. MpMRI) and with training set filtering at various rates. The evidential focal loss described in Section 4.2.2 is used to optimize the models. The co-teaching framework is excluded from this section for the following reason: while co-teaching implicitly handles noisy labels or samples in the training set, the training set filtering in Section 4.2.2 is an explicit alternative way of dealing with them. The co-teaching framework will first update its model parameters with simpler and cleaner samples during training. However, through the filtering process, data samples with high uncertainty values are considered potentially noisy and do not involve in the training process. We argue that it would be a duplicate procedure if we use co-teaching and filtering the training data simultaneously. Our hypothesis for training set filtering is by explicitly eliminating those highly uncertain data samples from the set and optimizing only on the rest of the “confident” samples using the evidential focal loss (Section 4.2.2), we could produce a more robust model. Therefore, to coalesce our proposed loss function with training set filtering, we do not use co-teaching and instead, we select MpMRI and M.S. MpMRI for experiments in this section. We use the MpMRI (resp. M.S. MpMRI) model to compute the uncertainty for all training data first, and then the filtering process can be either done in patch-driven or patient-driven on the training set, as we discussed in Section 4.2.2.
@@
In the second part of Table 1, we present the patient-based results on filtering 10% and 20% of training patches in the training set for the two selected models, while the third part of the table reports the results of filtering 10% and 20% of training patients. The results from both sections demonstrate that the binary classification performance improves when filtering more uncertain data for both models. Comparing these results with those from the first section of Table 1, we can conclude that the dataset filtering method applied to the training set, together with the evidential focal loss we proposed, can effectively improve the classification performance.
Moreover, an interesting observation is that the performance gradually deteriorates in patient-driven filtering. The reason behind this may be that in patch-driven filtering, we simply exclude some training patches with a high uncertainty value in the training process, no matter which patient the patches belong to. However, in the case of patient-driven filtering, we have to consider the average uncertainty of the 20 patches for each patient. If the average uncertainty of a patient is below the threshold, all the patches would be used for training, regardless of whether a specific patch has a very high uncertainty value. Therefore, there is a risk that we may falsely retain high uncertain patches because the corresponding patient has relatively low uncertainty on average, which can affect the model’s performance. This is also the reason why patch-driven filtration results are better.
@@
Ablation Study: As previously mentioned, the first section of Table 1 corresponds to experiments conducted without using evidential focal loss or filtering. On the other hand, the second and third sections encompass experiments that incorporate both these elements. In order to solely examine the influence of our proposed loss, we conducted an experiment where the evidential focal loss was employed without any filtering (0%). This results are summarized in Table 2 in comparison with 20% patch-based filtering approach. As can be seen, even without any data filtering during the training, we could correctly classify all patients with clinically significant PCa (sensitivity = 100%), which demonstrates a significant improvement compared to the baseline result in Grebenisan et al. (2021). As expected, the addition of data filtering further improves the results. The original results based on image patches of Table 2 can be found in Appendix A.
@@
Acc. Sen. Spec. AUC Acc. Sen. Spec. AUC
MpMRI 82.4 100 77.8 89.4$\pm$9.1 85.3 100 81.5 98.4$\pm$1.6
M.S. MpMRI 76.5 100 70.4 82.0$\pm$12.7 85.3 71.4 88.9 92.6$\pm$7.4
Table 2: Ablation on employing proposed evidential focal loss with and without data filtering for the two selected architectures. The Patient-based results are reported. Acc., Sen., Spec., and AUC are the shorts for Accuracy, Sensitivity, Specificity, and Area Under (ROC) Curve, respectively. Standard deviations are computed from the 95% bootstrap confidence interval with $n=3000$ samples. Best results are in bold. All units of the numeric values are in %.

### v2

@@ 4.2.2 Evidential Focal Loss
Patch-driven filtering: Given the uncertainty for each training patch, we simply eliminate $x\%,x\in[10,20]$ of the patches with the highest uncertainty and retrain the model on the rest of the samples on the training set.
@@ 5 Experiments & Setup
5.1 Setup
Data Split: For the domain transfer task, as mentioned in Section 3.2 and 4, the resampled T2 and ADC “images" with size $256\times 256$ from both ProstateX and the local dataset is used for training the ACL-GAN model. Particularly, we allocate 90% of images in both datasets as training and keep 10% as the validation set to avoid overfitting the ACL-GAN model. Importantly, we ensure that the images corresponding to each patient are exclusively present in either the training or validation set, but not both. To improve the robustness and enhance the ability of the ACL-GAN to capture feature-level representations of 1.5T images, we use all data from our local hospital. However, it is important to note that this approach does not yield any additional impacts on the subsequent classification task. The model only modifies image regions that have visual differences caused by acquisition parameters of various MRI machines, but it does not alter the context of the prostate itself.
For the PCa classification task, we use cropped and augmented T2 and ADC patches from both datasets. As mentioned before, this includes 204 ProstateX patients translated to 1.5T, as well as 104 patients from our local hospital captured in 1.5T. Regarding the data split for the classification, we keep patches of 34 patients from our local center as a standalone test set. From the remaining patches (70 local patients and all ProstateX patients), we allocate 80% for training and 20% for validation, assuring patches from the same patient are not included in both of these sets.
Next, we provide a brief intro to the experiments we conducted in this study.
Domain Transfer: The first experiment involved translating ProstateX MRI data from 3.0T to 1.5T using our proposed ACL-GAN model. We evaluated the effectiveness of our proposed approach by using quantitative metrics such as Fréchet Inception Distance (FID) score [103] and maximum mean discrepancy (MMD) score [104], more details could be found in Section 5.3. The translated MRI data was then employed in a downstream binary classification task for clinically significant PCa, demonstrating the superiority over the SDNet [5] baseline.
Classification: We divided our classification experiments into two folds. (1). Conventional approach, we used the conventional training paradigm without any filtering or uncertainty estimation. Three different model architectures were utilized for these experiments as discussed in Section 4.2.1. (2). Uncertainty-aware approach, we used the dataset filtering method and evidential focal loss proposed in Section 4.2.2 to train our models. Additionally, several ablation studies on data modalities, model architectures, and loss functions were conducted, with corresponding classification results detailed in Section 6. Finally, we focused on dataset filtering during deployment and examined how this technique affects the classification performance on the test data.
@@ 5.2 Experimental Details
We trained two ACL-GAN models separately for T2 and ADC images as part of our domain transfer framework. The optimizer used for both models was Stochastic Gradient Descent with Adam update rule [105], with an initial learning rate of 0.0001 and weight decay of 0.0001 to prevent overfitting. The batch size is 3 and are trained for 30,000 epochs. Moreover, when training the model for T2 images, we set the $\lambda_{mask}=0.0025,\lambda_{idt}=1,\lambda_{acl}=0.2$ in Equation (6) and lower and upper mask threshold to be 0.005 and 0.1, respectively. When training the model for ADC images, the values of $\lambda_{mask},\lambda_{idt},\lambda_{acl}$ are the same as in the T2 model with lower and upper mask thresholds set to 0.001 and 0.005, respectively. We adopt the Least-Square (LS) loss [106] for $\mathcal{L}_{adv}$ and $\mathcal{L}_{acl}$ in Equation (6), as done in [61].
Converting 3.0T to 1.5T: Once we have obtained two ACL-GAN models, we need to standardize the acquisition parameters of 3.0T prostate MRIs to match those of the 1.5T data in our local dataset. To achieve this, we divided the original 3.0T MRI into multiple 2D grayscale slices. For each 2D slice, we used the generator $G_{T}$ and a noise vector $z$ randomly sampled from $\mathcal{N}(0,1)$ to translate the 3.0T slice to 1.5T, e.g., $I_{\text{1.5T}}=G_{T}(I_{\text{3.0T}},z)$ as discussed in Section 4.1. This process was repeated for all 2D slices, and the slices were stacked back together to reconstruct the 3D MRI for each patient. The voxel spacing remained unchanged before and after the translation process. This procedure was applied to both T2 and ADC data.
All classification models were trained with Stochastic Gradient Descent with Adam, and batch normalization was applied to expedite convergence. In the first category of classification experiments (Conventional approach), the traditional focal loss [36] with $\gamma=2$ was employed. Specifically, all models except co-teaching were trained for 300 epochs with a learning rate of 0.0001, weight decay of 0.01, and a batch size of 10. For the co-teaching model, the noise rate and forget rate were set to 0.1, and the number of epochs for the linear drop rate to 10. The model was trained for 300 epochs, with a batch size set to 10, and a learning rate of 0.00001. In the second category of experiments (Uncertainty-aware approach), a learning rate of 0.0001 was used and decayed by a factor of 0.1 every 200 epochs. The weight decay was set to 0.01, the total training epochs were 300, and the batch size was 10. Further training details can be found in Appendix A.
@@ 5.3 Evaluation
To access the quality of the translated images and validate the effectiveness of our proposed domain transfer framework for ProstateX data from 3.0T to 1.5T, we followed previous works [78, 107, 108] to compute the maximum mean discrepancy (MMD) score [104] and the Fréchet Inception Distance (FID) score [103]. We followed the implementation provided in the Project-MONAI library  [^1] https://github.com/Project-MONAI/GenerativeModels/tree/main/generative/metrics. Both MMD and FID score measure the distribution distance between translated 1.5T images from the ProstateX dataset and real 1.5T images from our local hospital. Lower values in these metrics indicate greater fidelity to the real data distribution. The translated 1.5T images’ quality was further validated in a downstream classification task, combining them with the original 1.5T images to train a classifier distinguishing patients with and without clinically significant PCa. Traditional classification metrics, including accuracy (Acc.), sensitivity (Sen.), specificity (Spec.), and Area under ROC curve (AUC), were employed for this evaluation. Furthermore, we used “uncertainty calibration” to assess for performance of our uncertainty-based models. To compute calibration, we employed Expected Calibration Error (ECE) as done in [109, 110]. ECE measures the correspondence between the confidence of predictions and the actual model accuracy. Unlike other classification metrics, smaller values of ECE indicate less miscalibration, hence indicating better model calibration capability.
Reporting the patient-level performance is more relevant to the real clinical setting. However, since patches were used as model input, aggregating individual results from these patches is necessary to calculate patient-level metrics. This involved using the classifier to predict test patches, which were then sequentially grouped into $x$ groups, where $x$ is the number of patients in the test dataset. For each patient, there are 20 patches hence probabilities due to the mentioned augmentation in Section 3.2. The median probability $\tilde{p}{i}$ was computed over 20 probabilities as the aggregated probability for each patient. Finally, a threshold of 0.5 was applied to determine whether a patient has PCa, assigning label 1 if $\tilde{p}{i}>0.5$ and 0 otherwise.
@@ 6 Results and Discussion
In this section, we analyze the translated image from our proposed domain transfer framework and report patient-based classification results; the performance of our methods on patches are reported in Appendix B.
6.1 Quantitative Analysis of Translated Samples
We evaluate the quality of translated T2 images using two common metrics MMD [104] and FID [103] score, as reported in Table 1. The FID score is computed over each slice in the MRI data, i.e., a $160\times 160\times 32$ MRI volume results in 32 individual slices and FID scores, and we take the average scrore across all slices as the final score. The ACL-GAN model we used outperforms the SDNet in both metrics, we attribute the observed improvements to several key factors. Unlike SDNet [5], which merges modality features from randomly selected 1.5T images with anatomical features from 3.0T, our method learns the overall data distribution in an adversarial manner. This enables it to capture the entire distribution of 1.5T images and perform the translation more effectively. Additionally, our approach ensures that the translated image retains crucial features of the original, with the generator making modifications only to specific parts of the image. These designs contribute to the superior performance of our ACL-GAN-based domain transfer framework.
Since SDNet [5] is solely trained on T2 images, a direct quantitative comparison for translated ADC images is omitted. Instead, we conducted an ablation study to highlight the improvement in classification performance of adding translated ADC images to our model, see details in Section 6.2.
MMD $\downarrow$ FID $\downarrow$
SDNet [5] 0.0018 12.740
ACL-GAN (ours) 0.00054 11.464
Table 1: Quantitative results of translated T2 images using the baseline method SDNet and ACL-GAN in our proposed domain transfer framework. Lower values indicate better performance for all metrics. Bold values represent the best results.
@@ 6.2 PCa classification without filtering
Table 2 summarizes the experiment results in this section, which contains the PCa classification performance using the conventional approach, as detailed in Section 5.1. We observed that the AUC of using the co-teaching framework with MpMRI architecture as the base model achieves the best AUC and outperforms the baseline. The co-teaching model exhibits approximately a 50% increase in sensitivity while experiencing only a modest  10% decrease in specificity compared to the baseline model. This suggests that the co-teaching model demonstrates superior learning capabilities for classifying both positive and negative data samples on the test set. In the training process, we adopt a greedy approach of assuming 10% of the samples to be noisy. Consequently, both models need to designate a portion of the data in each batch as “clean" to update the parameters. This strategy allowed our model to prioritize learning from the clean data, leading to enhanced robustness.
Ablation Study: We embed the results of the ablation study in Table 2, which explores two key variations: alteration of the number of input modalities and alteration of the architecture of the model. To assess the impact of data modalities on classification performance, the T2-only model is compared with MpMRI and M.S. MpMRI models, both use T2 and ADC patches as input. The addition of the ADC modality leads to a substantial improvement in classification performance, highlighting the utility of multi-modal information in guiding the model for clinically significant PCa classification. The examination of model architecture reveals that the model with simpler inputs, MpMRI, performs better. Furthermore, the results can be further enhanced by leveraging the co-teaching framework.
Data Acc. Sen. Spec. AUC
SDNet[5](baseline) T2 79.4 28.6 92.6 76.2$\pm$17.5
T2-only T2 64.7 71.4 63.0 77.8$\pm$18.0
MpMRI T2+ADC 79.4 85.7 77.8 84.7$\pm$15.5
Vol. MpMRI T2+ADC 67.6 71.4 66.7 68.9$\pm$26.1
M.S. MpMRI T2+ADC 73.5 71.4 74.1 82.5$\pm$14.3
MpMRI+co-teaching T2+ADC 82.3 85.7 81.2 88.4$\pm$10.6
Table 2: Patient-based results of experiments using conventional training paradigm in Section 5.1. Standard deviations are computed from the 95% bootstrap confidence interval with $n=3000$ samples. The best results are bold. All units of the numeric values are in %.
@@ 6.3 PCa classification with filtration
In this section, we embark on experiments utilizing two different architectures, MpMRI and M.S. MpMRI, and incorporating training set filtering at various rates. The evidential focal loss described in Section 4.2.2 is used as the loss function to optimize the models. The co-teaching framework is excluded from this section for the following reason: while co-teaching implicitly handles noisy labels or samples in the training set, the training set filtering in Section 4.2.2 is an explicit alternative to dealing with them. The co-teaching framework will first update its model parameters with simpler and cleaner samples during training. However, through the filtering process, data samples with high uncertainty values are considered potentially noisy and hence are not involved in the training process. We argue that the use of co-teaching and simultaneous data filtering might be redundant. Our hypothesis for training set filtering is that by explicitly eliminating highly uncertain data samples from the set and optimizing only on the remaining "confident" samples using the evidential focal loss (Section 4.2.2), we can produce a more robust model. Therefore, to coalesce our proposed loss function with training set filtering, we do not use co-teaching and instead, we select MpMRI and M.S. MpMRI for experiments in this section. We use these two models to compute the uncertainty for all training data first, and then the filtering process can be done either patch-driven or patient-driven on the training set, as we discussed in Section 4.2.2.
@@
In Table 3, we present the patient-based results on filtering 10% and 20% of training patches in the training set for the two selected models. Table 4 represents the same as Table 3, except we filter 10% and 20% of training patients in the training set. The results from both tables demonstrate that the MpMRI model performs better than the M.S.MpMRI model in each filtration rate, and the binary classification performance improves when filtering more uncertain data for both models. From the expected calibration error (ECE), we observed that the MpMRI model with 20% filtering on both patch- and patient-based has a lower ECE value compared to those for the M.S. MpMRI model, demonstrating the predicted output probabilities of the MpMRI model matches well with the actual probabilities of the ground truth. Comparing these results with those from Table 2, we can conclude that the dataset filtering method applied to the training set, together with the evidential focal loss we proposed, can effectively improve the classification performance.
Data F.R. F.M. Acc. Sen. Spec. AUC ECE $\downarrow$
MpMRI T2+ADC 10% patch 82.4 85.7 81.5 85.7$\pm$9.9 0.27
M.S.MpMRI T2+ADC 10% patch 82.4 71.4 85.2 83.6$\pm$13.5 0.15
MpMRI T2+ADC 20% patch 85.3 100 81.5 98.4$\pm$1.6 0.20
M.S. MpMRI T2+ADC 20% patch 85.3 71.4 88.9 92.6$\pm$7.4 0.22
Table 3: Patient-based results of experiments using evidential focal loss and patch-based filtering. Standard deviations are computed from the 95% bootstrap confidence interval with $n=3000$ samples. The “F.R.” and “F.M.” represent the filtering rate and the filtering method, respectively. The best results for each filtration rate are bold. All units of the numeric values are in %.
Data F.R. F.M. Acc. Sen. Spec. AUC ECE $\downarrow$
MpMRI T2+ADC 10% patient 88.2 100 85.2 92.6$\pm$7.4 0.27
M.S. MpMRI T2+ADC 10% patient 85.3 100 81.5 86.2$\pm$9.0 0.24
MpMRI T2+ADC 20% patient 73.5 85.7 70.4 86.8$\pm$12.6 0.21
M.S. MpMRI T2+ADC 20% patient 73.5 71.4 74.1 84.6$\pm$14.2 0.22
Table 4: Patient-based results of experiments using evidential focal loss and patient-based filtering. Standard deviations are computed from the 95% bootstrap confidence interval with $n=3000$ samples. The “F.R.” and “F.M.” represent the filtering rate and the filtering method, respectively. The best results for each filtration rate are in bold. All units of the numeric values are in %.
Moreover, an interesting observation reveals a gradual deterioration in performance with patient-driven filtering. This phenomenon may stem from the distinction between patch-driven and patient-driven filtering approaches. In patch-driven filtering, the exclusion of training patches with high uncertainty values is straightforward during the training process, irrespective of the patient to which these patches belong. Conversely, in the case of patient-driven filtering, we have to consider the average uncertainty of the 20 patches for each patient. In cases where the average uncertainty for a patient falls below the defined threshold, all patches for that patient are used in training, irrespective of whether specific patches exhibit exceptionally high uncertainty. Consequently, there exists a risk of erroneously retaining patches with high uncertainty if the average uncertainty for the corresponding patient is relatively low, potentially impacting the model’s performance. This nuanced difference contributes to the observed discrepancy in results, favoring the patch-driven filtration approach.
@@
Ablation Study: As previously mentioned, Table 2 corresponds to experiments conducted without using evidential focal loss or filtering. On the other hand, Table 3 and 4 encompass experiments that incorporate both these elements. To solely examine the influence of our proposed loss, we conducted an experiment employing the evidential focal loss without any filtering (0%). These results are summarized in Table 5 in comparison with 20% patch-based filtering approach. As can be seen, even without any data filtering during the training, we could correctly classify all patients with clinically significant PCa (sensitivity = 100%), which demonstrates a significant improvement compared to MpMRI using traditional focal loss in Table 2 and also the baseline result in [5]. As expected, the addition of data filtering further improves the results for all metrics. The original results based on image patches of Table 5 can be found in Appendix B.
@@
Acc. Sen. Spec. AUC ECE Acc. Sen. Spec. AUC ECE
MpMRI 82.4 100 77.8 89.4$\pm$9.1 0.29 85.3 100 81.5 98.4$\pm$1.6 0.20
M.S. MpMRI 76.5 100 70.4 82.0$\pm$12.7 0.23 85.3 71.4 88.9 92.6$\pm$7.4 0.22
Table 5: Ablation on employing proposed evidential focal loss with and without data filtering for the two selected architectures. The Patient-based results are reported. Standard deviations are computed from the 95% bootstrap confidence interval with $n=3000$ samples. The best results for each filtration rate are in bold. All units of the numeric values are in %.